# Modern approaches to building effective interpretable models of the property market using machine learning

### v1

@@ Abstract
In this article, we review modern approaches to building interpretable models of property markets using machine learning on the base of mass valuation of property in the Primorye region, Russia. The researcher, lacking expertise in this topic, encounters numerous difficulties in the effort to build a good model. The main source of this is the huge difference between noisy real market data and ideal data which is very common in all types of tutorials on machine learning. This paper covers all stages of modeling: the collection of initial data, identification of outliers, the search and analysis of patterns in data, the formation and final choice of price factors, the building of the model, and the evaluation of its efficiency. For each stage, we highlight potential issues and describe sound methods for overcoming emerging difficulties on actual examples. We show that the combination of classical linear regression with interpolation methods of geostatistics allows to build an effective model for land parcels. For flats, when many objects are attributed to one spatial point the application of geostatistical methods is difficult. Therefore we suggest linear regression with automatic generation and selection of additional rules on the base of decision trees, so called the RuleFit method. Thus we show, that despite the strong restriction as the requirement of interpretability which is important in practical aspects, for example, legal matters, it is still possible to build effective models of real property markets.
@@ 1 Introduction
The problem of forecasting the real estate value has existed for a long time and is actual nowadays. It could be approached differently and include different constraints that impact predicted values. For example, it could be the aim to get the most accurate predictions; it could be the limit on algorithm runtime; it could be the goal to get the interpretable model, and so on. The problems faced by researchers are the insufficiency of data; the market of interest is undeveloped; explanatory factors are insufficient for prediction of the real estate value and so on. There are also problems with the nonlinear influence of some variables on others or the low correlation between explanatory variables and dependent variable. All of the above complicate the prediction process and make it nontrivial.
@@
The scope of research could be different – starting with the prediction of the price of one property unit up to a mass valuation of real estate in some geographic area. The last can be applied for research of the property market or administrative regulations such as property taxes. In most cases in the Russian Federation property taxes are determined from the cadastral value of the object which is calculated in the process of cadastral valuation. Cadastral valuation is a process of estimation of the cadastral value of property to create a clear and valid system of taxation and other relations governed by law. In the case of land parcels it could be a calculation of rental charge for land parcel use, the definition of land redemption value or buying from public domain or municipal property; for buildings and building spaces – calculation of taxes for buy/sell, rental and exchange of object, registration of succession rights, definition of exchange value, registration of gift agreement. The process of determination of cadastral value requires taking into account various factors, related to location, size, and quality of real estate property and also expert assessment of its potential for use in various purposes (Byrda, 2025). Cadastral valuation significantly influences the economic development of territories in municipalities. In the face of modern economic changes and financial difficulties, the importance of correct regulation of cadastral valuation as one of the factors stimulating the growth of municipal territories becomes particularly important. The precise cadastral valuation promotes the creation of a true picture of real market property values and leads to the improvement of the investment environment and stimulation of the outlays. Incorrect cadastral valuation or ineffective using results of cadastral valuation could reduce the investment attractiveness and give rise to a growth rate reduction of local economies and a decline of the social and economic situation. Overcharge of territories by unreasonably high values of cadastral value takes a toll on the interests of potential investors and may cause secular stagnation in the property market. Non-optimal instruments of cadastral valuation ignoring the physical-geographic factor of pricing instigates maldistribution of tax burden and as a result, the insufficient fundraising for covering municipalities’ needs. Real property tax based on valid cadastral valuation can be used for sponsorship of municipal projects focusing on the development of local amenities and the social sector. Incorrect cadastral valuation can prevent the development of particular areas and create financial and economic disbalance (Byrda, 2025).
@@
Cadastral valuation is conducted by methods verging its predictions to market values as close as possible. The market value of an object can vary significantly over time. Due to this reason, the cadastral value of an object should be regularly updated through mass cadastral valuation of real estate. The result of such mass valuation is the model based on market data which predicts the price of the arbitrary object. In this paper, we discuss the methods of effective prediction for mass cadastral valuation of real estate on the example of state cadastral valuation. Initially, it means that such valuation should obey current law. In Russia, it states that the valuation process is regulated by methodological rules of mass valuation of real estate. According to them, the prediction of real estate value can be made using one of the next methods – comparative method, cost method, or income method. The comparative method is typically used in the case when the market of real estate property is highly developed and there is enough information about deals and offers. This method includes method of fiducial object and models of regression analysis such as linear regression and its derivatives (Ministry for Economic Development of Russia, 2018). However, in the modern world, where technologies improve every day and machine learning actively penetrates all spheres of life, making predictions more and more precise, it becomes clear that we need more widely implement such methods in the mass valuation of real estate to improve the existing algorithms.
@@
Today there are many scientific works on the application of machine learning to prediction of property value. There are not only simple models such as linear regression and its variations (Adair et al., 1996) but also more complex ones, for example, Support Vector Machines (Phan, 2018) or a combination of several methods (Gu et al., 2011). In the review of mass valuation models (Jahanshiri et al., 2011), authors distinguish linear regression as a basic model and highlight its main limitation – it deals badly with spatial characteristics of data. It is important in the case of real estate property. For overcoming this problem authors suggest various spatial methods (Jahanshiri et al., 2011). But if models are created to form the base for taxing then transparency becomes the key aspect of the model. In this context transparency means interpretability. The results of the model should be presented in an accessible for ordinary person format – to explain clearly how input data influences on final results. In most cases, the machine learning models are black box – the programmer and users see the input data and the results but the mechanism of calculation is rather implicit. Appraiser can more deeply understand these calculation principles and algorithms, and what methods are used for the adjustment of estimates, but explaining it to an ordinary person who only wants to know the answer to just one question “How was the value of my property calculated?”, becomes almost impossible task. The interpretability of the model allows to find patterns and estimate how factors influence the price. It gives an opportunity to explain the constitutes of price to citizens who express disagreement with calculated cadastral values. Citizens can estimate the correctness of included factors and see their values for their property. In case of litigation, the presence of clear and sound cadastral values can help to protect it in law-court. Obviously, this argument is inapplicable in the case of black box models.
@@
Nowadays there are many reference materials on machine learning, but most of them demonstrate concepts on the base of ideal data. These data significantly differ from real data that is subject to errors and fluctuation. As a result, when a researcher tries to practice ideas and methods from a book, he is faced with difficulties – either methods don’t work at all or work incorrectly. For solving this problem a researcher should undertake careful preparatory work with data. Unfortunately, tutorials do not pay much attention to this problem.
@@
Also in this paper we present approaches for solving problems involving the identification of outliers, and the creation and selection of explanatory variables. Of note, we work with two market segments – flats and land parcels. For each segment, we analyzed the data and built the model. Sections 3 “Methods of identification of the outliers” and 4 “The formation and selection of explanatory variables” present the basic principles and interesting facts of the research performed for both segments. The section 5 “Modeling” consists of two parts describing each market segment. This section presents a detailed description of the methods and principles of modeling.
2 Description of target of research
@@
The modeling territorially covers Primorsky Krai – the administrative region of the Russian Federation located in the southeast of the country (Figure 1). The region borders China on the west, North Korea on the south-west, and Khabarovsk Krai (Russian Federation) on the north. Primorsky Krai is the part of Far Eastern Federal District. The area of the region is 164 673 km2 which is comparable to areas of such countries as Bulgaria (110 994 km2), Greece (131 957 km2) or Cambodia (181 035 km2). The population of the region is 1 806 393, the administrative capital is the port city Vladivostok (population 591 628). Vladivostok is the political, cultural, scientific, and educational center of the region. In 2012 Vladivostok was the venue for the APEC (Asia-Pacific Economic Cooperation) summit. For this event a wide range of significant measures was taken – the development of telecommunication, the construction of new roads, enhancing the cultural and prestigious status of the region (building of Opera and Ballet Theatre, building of multipurpose health center), the reconstruction of energy and heat supply facilities contributing to the reduction of energy costs of enterprises and energy losses in public utilities. Vladivostok is one of ten cities in the country where the Federal University was created (its new campus on Russky Island was the venue for APEC). This university has become a powerful intellectual base for the region.
@@
The membership of Russia in APEC and Primorsky Krai in particular as the main territory for interaction with other countries of the Asia-Pacific region allows balanced management of the regional economy and effective use of the natural resource potential of the Far East of Russia. This in turn increases the attractiveness of the region as a place for residence and creates opportunities for enhancing the business availability and for recruitment of potential investors (Chebotarev, 2012). Additionally, Asia-Pacific Economic Cooperation favors regional economic integration, supports a multilateral trading system, unlocks the potential of the digital economy, and helps to bridge the digital divides (Ministry for Economic Development of Russia, ).
@@
Precisely because of its favorable territorial location Vladivostok been in a third ten of Russia’s cities by population and separated by thousands of kilometers from the central part of Russia stably occupies a place in the top ten of cities of Russia by real property value. But the modeling of real property value is a difficult task due to several factors: complex terrain, historico-geographical urban development, various transport accessibility of different parts of the city, the existence of insular territories, the mixing of historic and modern buildings, nonuniform density of social, cultural and entertainment centers, coexistence of industrial estate and domestic development and so on. The requirement of interpretability of the model additionally complicates the task. Thus one should at each stage use the wide set of machine learning methods for building an effective model of property values for Vladivostok.
@@ 3 Identification of outliers
An important step in modeling is checking data for validity. The data is prone to mistakes of several origins. The most common of them are accidental mistakes (typos and so on) and intentional distortion of data by publishing fictitious values – too high or too low. As a result, the process of creating a model becomes more difficult and may lead to incorrect results. It’s worth noting that in the process of analyzing the data on the outliers, one should take into account the type of objects. A land parcel is basically a one-dimensional object. Usually, the price of such an object depends on its distance relative to other objects, for example center of the town. On the other hand, the value of a flat depends not only on its location but also on the characteristics of the multi-unit residential building where the flat is located. In the second case, the process of identification of outliers can be more difficult. In the case when we analyze the one-dimensional objects (land parcels), we can plot their per square meter price on a map and estimate their validity from general considerations.
@@
Figure 2 shows that the points located close to each other can have significantly different per square meter prices. This may be caused by individual characteristics of the object – the land parcel could be in a zone with a high risk of flooding or the land parcel could have a significant slope. All these factors make land parcel exploitation more difficult. However, the determination of the cadastral value of the object is the mass process reflecting the value of a typical object without delving too much into its specific characteristics. Because of this, reason we can expect a similar per square meter price among land parcels in some local area that changes with distances to other objects included in price factors (center of a town, water objects, social objects, etc.). For the exclusion of outliers, we can start with the clusterization of points on the base of their spatial relationship. The next thing we can calculate the median or average value for each cluster and find the outliers using such methods as box-plot or z-score or so on. Also, such clustering helps to take into account the inhomogeneity of administrative territory. For example, the territory could include the city and the suburbs where a per square meter price is sufficiently different (Figure 3). On the contrary, if we consider the whole territory without clusterization we can lose part of the data because firstly the per square meter price of the city is much higher than the per square meter price of the suburbs and secondly the number of data points in the city are much less than the corresponding number in suburbs. In this case, the land parcels located in the city would be outliers and by excluding them we will have major losses of important data and our predicted value would be sufficiently distorted.
@@
Also, we can find the outliers using the probability density function of dependent variable. For example, Figure 4 shows that values between 50000 rub and 100000 rub for the deals are doubtful.
@@
We can confirm this assumption by taking buildings and analyzing per square meter prices in each building. The example of one building is shown in Table 1 where suspicious value is highlighted. Such deals could be fictitious and if we wouldn’t exclude them they would have a negative effect on the forecasting.
@@
offer | 5 | 3 850 000 | 17,1 | 225 146 |
offer | 2 | 4 500 000 | 23,1 | 194 805 |
offer | 8 | 6 150 000 | 34,5 | 178 261 |
offer | 9 | 4 350 000 | 17,0 | 255 882 |
offer | 3 | 4 200 000 | 17,0 | 247 059 |
offer | 4 | 3 800 000 | 16,7 | 227 545 |
offer | 5 | 3 800 000 | 17,0 | 223 529 |
deal | 5 | 4 300 000 | 16,7 | 257 485 |
deal | 2 | 4 100 000 | 23,0 | 178 261 |
deal | 4 | 4 050 000 | 16,9 | 239 645 |
deal | 4 | 900 000 | 16,8 | 53 571 |
deal | 4 | 4 900 000 | 23,3 | 210 300 |
deal | 4 | 3 800 000 | 16,7 | 227 545 |
@@
More serious methods such as machine learning also can be used for the identification of outliers. Particularly it would help for many-dimensional objects e.g. flats. Firstly we should perform territorial clustering of points that results in the division of the populated place into regions (clusters). Next, we create the new clusters in each region on the base of the characteristics of the buildings e.g. “Wall material”, “Number of stories” and “Year built”. This procedure allows to highlight typical buildings. We can identify such clusters either manually on the basis of our expertise and knowledge of historic periods of city development or we can use machine learning methods implementing clustering algorithms. Such algorithms take as input the buildings in the space of the above-mentioned characteristics and return the class markers for each building. In this paper, we use the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm (Ester et al., 1996). This algorithm is based on the idea of closeness between points in the space of features. It means that DBSCAN groups the close points and marks as outliers the single points. The minimal number of points in a neighborhood for a point to be considered as a core point and the maximum distance between two points for one to be considered as in the neighborhood of the other are adjustable parameters of the algorithm (Moreira et al., 2005). By changing the first we can control the density of resulting clusters while the second parameter is the most important one and should be chosen accurately depending on the data and the metric. The metric specifies how the distances in the space of features are calculated. After the process of finding typical buildings is finished one can start the identification of outliers within each group. During the analysis, we noticed that there is a negative dependence between the per square meter price and the area of the flat. In addition, this dependence is not linear but rather exponential (Figure 5). The highlighting of the trend for such data is a nontrivial task. If there are signs of linear dependence then we can try RANdom Sample Consensus (RANSAC) regression (Kumar, 2020; Fischler and Bolles, 1981) – linear regression robust to outliers. This algorithm is an iterative procedure where every step assumes the forming of the sub-sample and fitting linear model. Then this model is applied to the total sample and residuals are calculated. Everything beyond the residual threshold is marked as outliers. Then the model retrains on other values and all steps are repeated. The new model and previous model are compared on the number of outliers. A model that contains a smaller number of outliers is preferable. In our case, we can transform the exponential dependence to linear by taking the natural logarithm of data. So for each cluster found by DBSCAN, we apply RANSAC regression to find the outliers and to highlight the trend. The dependent variable is the natural logarithm of the per square meter price and the explanatory variable is the area of flat. This procedure allows to find the trend for data with high fluctuations and noise and hinges on the reasonable assumption that flats in buildings with similar characteristics located within some compact territory should have similar per square meter prices. The points that are highly different from the trend are considered to be outliers. These values can be explained by the individual characteristics of the flat and can’t account for the mass valuation. The example of the results of RANSAC regression is shown in Figure 5, the points marked as outliers are not used for forecasting.

### v2

# Modern approaches to building interpretable models of the property market using machine learning on the base of mass cadastral valuation

@@ Abstract
In this article, we review modern approaches to building interpretable models of property markets using machine learning on the base of mass valuation of property in the Primorye region, Russia. The researcher, lacking expertise in this topic, encounters numerous difficulties in the effort to build a good model. The main source of this is the huge difference between noisy real market data and ideal data which is very common in all types of tutorials on machine learning. This paper covers all stages of modeling: the collection of initial data, identification of outliers, the search and analysis of patterns in the data, the formation and final choice of price factors, the building of the model, and the evaluation of its efficiency. For each stage, we highlight potential issues and describe sound methods for overcoming emerging difficulties on actual examples. We show that the combination of classical linear regression with interpolation methods of geostatistics allows to build an effective model for land parcels. For flats, when many objects are attributed to one spatial point the application of geostatistical methods is difficult. Therefore we suggest linear regression with automatic generation and selection of additional rules on the base of decision trees, so called the RuleFit method. Thus we show, that despite such a strong restriction as the requirement of interpretability which is important in practical aspects, for example, legal matters, it is still possible to build effective models of real property markets.
@@ 1 Introduction
The problem of forecasting the real estate value has existed for a long time and is actual nowadays. It could be approached differently and include different constraints that impact predicted values. For example, it could be the aim to get the most accurate predictions; it could be the limit on algorithm runtime; it could be the goal to get the interpretable model, and so on. The problems faced by researchers are the insufficiency of data; the market of interest is undeveloped; explanatory factors are insufficient for prediction of the real estate value and so on. There are also problems with the nonlinear influence of some variables on others or the low correlation between explanatory variables and a dependent variable. All of the above complicate the prediction process and make it nontrivial.
@@
The scope of research could be different – starting with the prediction of the price of one property unit up to a mass valuation of real estate in some geographic area. The last can be applied for research of the property market or administrative regulations such as property taxes. In most cases in the Russian Federation property taxes are determined from the cadastral value of the object which is calculated in the process of cadastral valuation. Cadastral valuation is a process of estimation of the cadastral value of property to create a clear and valid system of taxation and other relations governed by law. In the case of land parcels it could be a calculation of rental charge for land parcel use, the definition of land redemption value or buying from public domain or municipal property; for buildings and building spaces – calculation of taxes for buy/sell, rent and exchange of an object, registration of succession rights, definition of exchange value, registration of gift agreement. The process of determination of cadastral value requires taking into account various factors, related to location, size, and quality of real estate property and also expert assessment of its potential for use in various purposes (Byrda, 2025). Cadastral valuation significantly influences the economic development of territories in municipalities. In the face of modern economic changes and financial difficulties, the importance of correct regulation of cadastral valuation as one of the factors stimulating the growth of municipal territories becomes particularly important. The precise cadastral valuation promotes the creation of a true picture of real market property values and leads to the improvement of the investment environment and stimulation of the outlays. Incorrect cadastral valuation or ineffective usage of results of cadastral valuation could reduce the investment attractiveness and give rise to a growth rate reduction of local economies and a decline of the social and economic situation. Overcharge of territories by unreasonably high cadastral values takes a toll on the interests of potential investors and may cause secular stagnation in the property market. Non-optimal instruments of cadastral valuation ignoring the physical-geographic factor of pricing instigates maldistribution of tax burden and as a result, the insufficient fundraising for covering municipalities’ needs. Real property tax based on valid cadastral valuation can be used for sponsorship of municipal projects focusing on the development of local amenities and the social sector. Incorrect cadastral valuation can prevent the development of particular areas and create financial and economic disbalance (Byrda, 2025).
@@
Cadastral valuation is conducted by methods verging its predictions to market values as close as possible. The market value of an object can vary significantly over time. Due to this reason, the cadastral value of an object should be regularly updated through mass cadastral valuation of real estate. The result of such mass valuation is the model based on market data which predicts the price of the arbitrary object. In this paper, we discuss the methods of effective prediction for mass cadastral valuation of real estate on the example of state cadastral valuation. Initially, it means that such valuation should obey current law. In Russia, it states that the valuation process is regulated by methodological rules of mass valuation of real estate. According to them, the prediction of real estate value can be made using one of the next methods – comparative method, cost method, or income method. The comparative method is typically used in the case when the market of real estate property is highly developed and there is enough information about deals and offers. This method includes the method of a fiducial object and models of regression analysis such as linear regression and its derivatives (Ministry for Economic Development of Russia, 2018). However, in the modern world, where technologies improve every day and machine learning actively penetrates all spheres of life, making predictions more and more precise, it becomes clear that machine learning algorithms should be more widely implemented in the mass valuation of real estate to improve the existing methods.
@@
Today there are many scientific works on the application of machine learning to prediction of property value. There are not only simple models such as linear regression and its variations (Adair et al., 1996) but also more complex ones, for example, Support Vector Machines (Phan, 2018) or a combination of several methods (Gu et al., 2011). In the review of mass valuation models (Jahanshiri et al., 2011), authors distinguish linear regression as a basic model and highlight its main limitation – it deals badly with spatial characteristics of data. It is important in the case of real estate property. For overcoming this problem authors suggest various spatial methods (Jahanshiri et al., 2011). But if the model is created to form the base for taxing then transparency becomes its key aspect. In this context transparency means interpretability. The results of the model should be presented in an accessible for an ordinary person format – to explain clearly how input data influences on final results. In most cases, the machine learning models are black box – the programmer and users see the input data and the results but the mechanism of calculation is rather implicit. Appraiser can understand these calculation principles and algorithms, and what methods are used for the adjustment of estimates more deeply, but explaining it to an ordinary person who only wants to know the answer to just one question “How was the value of my property calculated?”, becomes almost impossible task. The interpretability of the model allows to find patterns and to estimate how factors influence the price. It gives an opportunity to explain the constitutes of price to citizens who express disagreement with calculated cadastral values. Citizens can estimate the correctness of included factors and see their values for their property. In case of litigation, the presence of clear and sound cadastral values can help to protect it in law-court. Obviously, this argument is inapplicable in the case of black box models.
@@
Nowadays there are many reference materials on machine learning, but most of them demonstrate concepts on the base of ideal data. These data significantly differ from real data that is subject to errors and fluctuations. As a result, the implementation of the ideas and methods from a book becomes challenging - they either do not work at all or work incorrectly. For solving this problem a researcher should undertake careful preparatory work with data. Unfortunately, tutorials do not pay much attention to this problem.
@@
Also in this paper we present approaches for solving problems involving the identification of outliers, and the creation and selection of explanatory variables. Of note, we work with two market segments – flats and land parcels. For each segment, we analyzed the data and built the model. In section 2 “Description of the target of research” we briefly introduce the territory for which we build the models and shortly describe the data for modeling. More detailed analysis of data is given in the following sections. Sections 3 “Identification of outliers” and 4 “Formation and selection of explanatory variables” present the basic principles and interesting facts of the research performed for both segments. The section 5 “Modeling” consists of two parts describing each market segment. This section presents a detailed description of the methods and principles of modeling. In section 6 “Conclusion” we outline the research and highlight the most important points.
2 Description of the target of research
@@
The modeling territorially covers Primorsky Krai – the administrative region of the Russian Federation located in the southeast of the country (Figure 1). The region borders China on the west, North Korea on the south-west, and Khabarovsk Krai (Russian Federation) on the north. Primorsky Krai is the part of the Far Eastern Federal District. The area of the region is 164 673 km2 which is comparable to areas of such countries as Bulgaria (110 994 km2), Greece (131 957 km2) or Cambodia (181 035 km2). The population of the region is 1 806 393, the administrative capital is the port city Vladivostok (population 591 628). Vladivostok is the political, cultural, scientific, and educational center of the region. In 2012 Vladivostok was the venue for the APEC (Asia-Pacific Economic Cooperation) summit. For this event a wide range of significant measures was taken – the development of telecommunication, the construction of new roads, the enhancement of the cultural and prestigious status of the region (building of the Opera and Ballet Theater, building of multipurpose health center), the reconstruction of energy and heat supply facilities contributing to the reduction of energy costs of enterprises and energy losses in public utilities. Vladivostok is one of ten cities in the country where the Federal University was created (its new campus on the Russky Island was the venue for APEC). This university has become a powerful intellectual base for the region.
@@
The membership of Russia in APEC and Primorsky Krai in particular as the main territory for interaction with other countries of the Asia-Pacific region allows balanced management of the regional economy and effective use of the natural resource potential of the Far East of Russia. This in turn increases the attractiveness of the region as a place for residence and creates opportunities for enhancement of the business availability and for recruitment of potential investors (Chebotarev, 2012). Additionally, Asia-Pacific Economic Cooperation favors regional economic integration, supports a multilateral trading system, unlocks the potential of the digital economy, and helps to bridge the digital divides (Ministry for Economic Development of Russia, ).
@@
Precisely because of its favorable territorial location Vladivostok being in the third ten of Russia’s cities by population and separated by thousands of kilometers from the central part of Russia stably occupies a place in the top ten of cities of Russia by real property value. But the modeling of real property value is a difficult task due to several factors: complex terrain, historico-geographical urban development, various transport accessibility of different parts of the city, the existence of insular territories, the mixing of historic and modern buildings, nonuniform density of social, cultural and entertainment centers, coexistence of industrial estate and domestic development and so on. The requirement of interpretability of the model additionally complicates the task. Thus one should at each stage use the wide set of machine learning methods for building an effective model of property values for Vladivostok.
@@
In practice, analysts prefer to model the PSMP but not the total estate price. The reason is that the “Area” is fairly significant pricing variable and non-inclusion it to the model will often result in decrease of quality of the model. But if one considers the total value of estate price and simultaneously includes the “Area” into a list of independent variables then it will just completely suppress other variables with no contribution to the predictive power of the model.
@@ 3 Identification of outliers
An important step in modeling is checking data for validity. The data is prone to mistakes of several origins. The most common of them are accidental mistakes (typos and so on) and intentional distortion of data by publishing fictitious values – too high or too low. As a result, the process of creating a model becomes more difficult and may lead to incorrect results. It’s worth noting that in the process of analyzing the data on the outliers, one should take into account the type of objects. A land parcel loosely could be considered as an one-dimensional object. We mean that basically the geographic coordinates of a land parcel define its price. Coordinates decode both the spatial location of the object and its individual characteristics such as the slope, the availability of electricity, and so on. There is one to one mapping between coordinates and the land parcel. On the other hand, the typical multi-unit residential building contains several tens or even hundreds of flats which means that one spatial point corresponds to multiple objects. Flats in the same building could differ by the storey or the area. Even the adjacent buildings having almost the same coordinates could differ drastically by the year of built, the wall material, etc. This is fairly common situation for flats and rather untypical for land parcels where the neighboring objects tend to have similar characteristics. As the result, the process of identification of outliers for the former could be more difficult. For the latter case, we can plot the per square meter price of land parcels on a map and estimate its validity from general considerations.
@@
Figure 2 shows that the points located close to each other can have significantly different per square meter prices. This may be caused by individual characteristics of the object – the land parcel could be in a zone with a high risk of flooding or the land parcel could have a significant slope. All these factors make land parcel exploitation more difficult. However, more likely, a such difference in prices between neighboring land parcels means that the one or both prices are anomalous and should be considered as outliers. The price of objects like land parcels tends to change rather continuously with distance between them without sharp variations. Because of this reason, we can expect a similar per square meter price among land parcels in some local area that changes with distances to other objects included in price factors (center of a town, water objects, social objects, etc.).
For the exclusion of outliers, we can start with the clusterization of points on the base of their spatial relationship. Next we can calculate the median or average value for each cluster and find the outliers using such methods as box-plot or z-score and so on. Also, such clustering helps to take into account the inhomogeneity of administrative territory. For example, the territory could include the city and the suburbs where a per square meter price is sufficiently different (Figure 3). On the contrary, if we consider the whole territory without clusterization we can lose part of the data because firstly the per square meter price of the city is much higher than the per square meter price of the suburbs and secondly the number of data points in the city is much less than the corresponding number in suburbs. In this case, the land parcels located in the city would be outliers and by excluding them we will have major losses of important data and our predicted value would be sufficiently distorted.
@@
Also, we can find the outliers using the probability density function of the dependent variable. For example, Figure 4 shows that values between 50000 rub and 100000 rub for the deals are doubtful.
@@
We can check this assumption by taking buildings and analyzing per square meter prices in each building. The example of one building is shown in Table 1 where suspicious value is highlighted. Such deals could be fictitious and if we wouldn’t exclude them they would have a negative effect on the forecasting.
@@
offer | 5 | 3 850 000 | 17.1 | 225 146 |
offer | 2 | 4 500 000 | 23.1 | 194 805 |
offer | 8 | 6 150 000 | 34.5 | 178 261 |
offer | 9 | 4 350 000 | 17.0 | 255 882 |
offer | 3 | 4 200 000 | 17.0 | 247 059 |
offer | 4 | 3 800 000 | 16.7 | 227 545 |
offer | 5 | 3 800 000 | 17.0 | 223 529 |
deal | 5 | 4 300 000 | 16.7 | 257 485 |
deal | 2 | 4 100 000 | 23.0 | 178 261 |
deal | 4 | 4 050 000 | 16.9 | 239 645 |
deal | 4 | 900 000 | 16.8 | 53 571 |
deal | 4 | 4 900 000 | 23.3 | 210 300 |
deal | 4 | 3 800 000 | 16.7 | 227 545 |
@@
More serious methods such as machine learning also can be used for the identification of outliers. Particularly it would help for “multi-dimensional” objects like flats. Firstly we should perform territorial clustering of points that results in the division of the populated place into regions (clusters). Next, we create the new clusters in each region on the base of the characteristics of the buildings e.g. “Wall material”, “Number of storeys” and “Year built”. This procedure allows to highlight typical buildings. We can identify such clusters either manually on the basis of our expertise and knowledge of historic periods of city development or we can use machine learning methods implementing clustering algorithms. Such algorithms take as input the buildings in the space of the above-mentioned characteristics and return the class markers for each building. In this paper, we use the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm (Ester et al., 1996) with realization on Python (the library “Scikit-learn” (Pedregosa et al., 2011)). This algorithm is based on the idea of closeness between points in the space of features. It means that DBSCAN groups the close points and marks as outliers the single points. The minimal number of points in a neighborhood for a point to be considered as a core point and the maximum distance between two points for one to be considered as in the neighborhood of the other are adjustable parameters of the algorithm (Moreira et al., 2005). By changing the first we can control the density of resulting clusters while the second parameter is the most important one and should be chosen accurately depending on the data and the metric. The metric specifies how the distances in the space of features are calculated. After the process of finding typical buildings is finished one can start the identification of outliers within each group. During the analysis, we noticed that there is a negative dependence between the per square meter price and the area of the flat. In addition, this dependence is not linear but rather exponential (Figure 5). The highlighting of the trend for such data is a nontrivial task. If there are signs of linear dependence then we can try RANdom Sample Consensus (RANSAC) regression (Kumar, 2020; Fischler and Bolles, 1981) – linear regression robust to outliers. This algorithm is an iterative procedure where every step assumes the forming of the sub-sample and fitting linear model. Then this model is applied to the total sample and residuals are calculated. Everything beyond the residual threshold is marked as outliers. Then the model retrains on other values and all steps are repeated. The new model and previous model are compared on the number of outliers. A model that contains a smaller number of outliers is preferable. In our case, we can transform the exponential dependence to linear by taking the natural logarithm of the data. So for each cluster found by DBSCAN, we apply RANSAC regression to find the outliers and to highlight the trend. The dependent variable is the natural logarithm of the per square meter price and the explanatory variable is the area of a flat. This procedure allows to find the trend for data with high fluctuations and noise and hinges on the reasonable assumption that flats in buildings with similar characteristics located within some compact territory should have similar per square meter prices. The points that are highly different from the trend are considered to be outliers. These values can be explained by the individual characteristics of the flat and can’t account for the mass valuation. The example of the results of RANSAC regression is shown in Figure 5, the points marked as outliers are not used for forecasting. We used median absolute deviation (MAD) as the threshold. The “SciKit-learn” library uses it by default.